OpenStudy (anonymous):

Geometry help please, What is the length of the missing side of the right triangle shown below?

OpenStudy (anonymous):

i hope it is a right triangle

OpenStudy (anonymous):

use \(a^2+b^2=c^2\) to find which ever side is missing

OpenStudy (anonymous):

ANSWER CHOICES: (4, 16, \[\sqrt{15}\] , \[\sqrt{34}\] )

OpenStudy (anonymous):

start with \(3^2+5^2=c^2\) and then compute